OpenStudy (idku):

Math - Based - Physics .... QUESTIONS. (I am working the parts 1-6 on my own, if you want, please come and view and correct if anything) A soccer ball is kicked with an initial horizontal velocity of 11 m/s and an initial vertical velocity of 17 m/s.

OpenStudy (feemeegaming):

AND?

OpenStudy (idku):

I got disconnected I apologize. I will post parts of the problem will do them here on my own and if I need help i will look for it.

OpenStudy (idku):

http://i.imgur.com/I0prCLo.png here is the IMAGE ------------------------------------------------------------ 1) What is the initial speed of the ball? ------------------------------------------------------------ 2) What is the initial angle θ of the ball with respect to the ground? ------------------------------------------------------------ 3) What is the maximum height the ball goes above the ground? ------------------------------------------------------------ 4) How far from where it was kicked will the ball land? ------------------------------------------------------------ 5) What is the speed of the ball 2.5 seconds after it was kicked? ------------------------------------------------------------ 6) How high above the ground is the ball 2.5 seconds after it is kicked? ------------------------------------------------------------

OpenStudy (idku):

` PART 1: ` The initial speed of the ball (lets denote it with u) is: \(\large u=\sqrt{11^2+17^2}=20.24\color{blue}{\left({\rm m/s}\right)}\)

OpenStudy (idku):

` PART 2:` \(\large \theta={\rm Tan}^{-1}\left({\rm 17/11}\right)\approx 57.09\)

OpenStudy (feemeegaming):

Sorry i dont know.

OpenStudy (idku):

Now for part 3, we know that maximum height is reached when v=0. Also, I am given: a = -9.81m/s\(^2\) u = 20.24m/s So will I use? \(v^2=u^2+2ax\) NOTE: "u" denotes initial speed. "x" denotes displacement/distance (in this case same thing) \(v^2=u^2+2ax\) \(0^2=17^2+2(-9.81)x\) \(-17^2=2(-9.81)x\) \(17^2=2(9.81)x\) \(x=17^2/[2(9.81)]\approx 14.73\)

OpenStudy (idku):

`PART 4` How far from where it was kicked will the ball land? `(Horizontal Displacement) = (Time) × (Horizontal Velocity)` I know the horizontal velocity, but the time is unknown. We can use the vertical components to find the time!! Since the ball travels in a parabolic motion it is just twice the time from v=17 (initial horiz velocity) to v=0 (the maximum height point, or the vertex) so we went from 17 m/s to 0 m/s how many second would that have took if every second we lose 9.81m/s? 17(m/s) / 9.81(m/s^2) ≈ 1.732 seconds Now remember we go twice from v=17m/s to v=0m/s So we really have 3.464 seconds of FULL motion time when ball is in air. units work out too! Now we can find the time. `(Horizontal Displacement) = (Time) × (Horizontal Velocity)` `(Horizontal Displacement) = (3.464) × (11)` `(Horizontal Displacement) = 38.104`

OpenStudy (idku):

I am having trouble with part 5.

OpenStudy (idku):

A soccer ball is kicked with an initial horizontal velocity of 11 m/s and an initial vertical velocity of 17 m/s. 5) What is the speed of the ball 2.5 seconds after it was kicked?

OpenStudy (idku):

v=u+at u - initial speed t - time a - accel due to gravity = -9.81m/s^2 v=20.24+(-9.81)2.5=-4.285 speed is always positive, so 4.285 because we want to know what happens with speed at t=2.5 No, that doesn;t work for some reason.

OpenStudy (anonymous):



OpenStudy (idku):

11+(9.81)2.5=35.525 and that is incorrect either (I put it in my practice thingy and it denied)

OpenStudy (idku):

oh, 20.24 not 11, I suppose. initial velocity V< not speed of horiz.

OpenStudy (idku):

20.24+(9.81)2.5=44.675 but that is too large. and I check it is not right

OpenStudy (anonymous):